3. You are given the following exchange rates: Exchange Rates Time 0 Time 1 So Si USD / GBP 1.8558 1.8561 USD/EUR 1.2674 1.26g. What is the AUD/ZAR cross rate? h. What is the ZAR / GBP cross rate? i. What is the EUR/GBP cross rate?
Answer #1

a) GBP AND EUR

b) CAD, MXN, AUD and ZAR

To calculate percentage change in a particular currency, you will first need to make it the base currency and then calculate.

c) No conversion required since GBP is already the base currency

% change =(1.8561/1.8558)-1

= 0.0162%

d) As AUD is not the base currency, we will need to convert it. Take reciprocal to convert to the base currency

USD/AUD (at time 0) = 1/1.3095

= 0.7637

USD/AUD (at time 1) = 1/1.3025

= 0.7678

% change = (0.7678/ 0.7637)-1

= 0.5369%

e) % change = (1.2622/1.2674)-1

= -0.4103%

f) take reciprocal to convert

USD/MXN (at time 0) = 1/10.7575

= 0.0930

USD/MXN (at time 1) = 1/10.6780

= 0.0937

% change = 0.7527%

g) At time 1, AUD/ZAR = AUD/USD * (1/ZAR/USD)

= (1.3025*(1/6.8850))

= 0.1892

h) ZAR/GBP = (ZAR/USD)*(USD/GBP)

= 12.7236

I) EUR/GBP= (1/(USD/EUR))* (USD/GBP)

= 1.4705